Calculate the weighted average cost of capital for the following firm: it has $600000 in debt, $400000 in common stock and $200000 in preferred stock. It has a 4% cost of debt, 14% cost of common stock, 12% cost of preferred stock and a 34% tax rate.

Group of answer choices

7.99%

8.64%

5.14%

9.3%
Answer #1

After-tax cost of debt=4*(1-tax rate)

=4*(1-0.34)=2.64%

Total value=(600000+400000+200000)=$1200000

WACC=Respective costs*Respective weight

=(600000/1200000*2.64)+(400000/1200000*14)+(200000/1200000*12)

=7.99%(Approx).